OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x – 4| + 1 = 8 Step 2: |x – 4| = 8 – 1 Step 3: |x – 4| = 7 Which of the following is a correct next step to solve the equation? x – 4 = –7 x + 4 = –7 –x – 4 = 7 x + 4 = 7

OpenStudy (anonymous):

I am thinking it is the first answer choice

OpenStudy (anonymous):

what about this one? Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) greater than or equal to -31? x greater than or equal to -9 -9 less than or greater to x A number line with a closed circle on -9 and shading to the left A number line with a closed circle on -9 and shading to the right

OpenStudy (anonymous):

8x - (5x + 4) >= -31 8x -5x - 4 >= -31 3x - 4 >= -31 3x >= -31 + 4 3x >= - 27 x >= -9 the incorrect one is : closed circle on -9 shading to the LEFT

OpenStudy (anonymous):

thank you... so much! do you ,mind helping with a couple more?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 3 < -9 or x + 5 greater than or equal to 10. A number line with an open circle on -6, shading to the left, and a closed circle on 5, shading to the right. A number line with a closed circle on -6, shading to the left, and an open circle on 5, shading to the right. A number line with an open circle on -6, a closed circle on 5, and shading in between. A number line with a closed circle on -6, an open circle on 5, and shading in between.

OpenStudy (anonymous):

x - 3 < -9 x < -9 + 3 x < - 6 x + 5 >= 10 x >= 10 - 5 x >= 5 open circle on -6, shaded to the left, and a closed circle on 5, shading to the right

OpenStudy (anonymous):

quick lesson... if there is an equal sign involved, it is a closed circle if there is no equal sign, it is an open circle (>) greater then means shading to the right (<) less then means shading to the left

OpenStudy (anonymous):

I am sorry but I have to go now.....I hope I helped you out some :)